Abstract. We evaluate the determinant

det_{0≤i,j≤n−1} (δ_{ij} + ∑_{t,k=0}^{n−1} \binom{i + μ}{t}\binom{k + ν}{k − t}\binom{j − k + μ − 1}{j − k}2^{k−t})

which gives the 2-enumeration of certain shifted plane partitions. This generalizes a result of Andrews (Aequationes Math. 33 (1987), 230– 250), who evaluated this determinant for ν = 0, thereby proving a conjecture of Mills, Robbins and Rumsey (Discrete Math. 67 (1987), 43–55).